### v2

@@ 4.1. RQ1: What Security DSLs have been presented in the scientific literature?
Most papers (49 %) have been published on venues targeting the security community, followed by 11 % of venues in software engineering. Networks (9 %), applied computing (7 %), information systems (7 %), programming languages (7 %), and distributed computing (6 %) venues are other commonly observed communities. Note that a single venue may target multiple communities. A significant portion of the venues (3 %), we could not assign to specific communities because they are general in nature and do not target identifiable communities, but rather computer scientists in general. For example, this journal falls into the category of publication venues targeting computer scientists in general. In Figure 2(b), we list all conferences and journals where at least four of the papers we examined were published, as well as the communities to which we assigned them.
@@
0010101010202020203030303040404040505050506060606070707070SecuritySoftware EngineeringNetworkProgramming LanguagesInformation SystemsApplied ComputingDistributed ComputingTheoretical Computer ScienceOperating SystemsHardwareRequirementsArtificial IntelligenceOtherNumber of Papers
@@
Only 40 of the 120 identified security DSLs (33 %) have a publicly accessible repository or website. Note that our initial search for publications covers the time span between 2010 and 2022 and with the backwards snowballing we additionally identified relevant papers dating back until 1997. Many security DSLs have been updated even after their release, as shown in Fig. 3. In particular, 25 % of the updates happened in 2023 or the first half of 2024, indicating that quite a few of them are actively maintained and possibly also used in practice or academia.
19971997199719971998199819981998199919991999199920002000200020002001200120012001200220022002200220032003200320032004200420042004200520052005200520062006200620062007200720072007200820082008200820092009200920092010201020102010201120112011201120122012201220122013201320132013201420142014201420152015201520152016201620162016201720172017201720182018201820182019201920192019202020202020202020212021202120212022202220222022202320232023202320242024202420240022224444666688881010101012121212Number of DSLspublishedlast update
@@
Observation 1. Security DSLs were primarily published in the mid-2010s, mainly from the security community, but also from other communities such as software engineering. While few DSLs are publicly available, many of those that are are still actively maintained.
@@ 4.2. RQ2: What security aspects are targeted by the security DSLs?
In this work we only cover DSLs that are security specific. To get an overview over the security aspects covered by the DSLs we analyze their fundamental security objective and provide a coarse-grained categorization according to the security concept and the system scope targeted by the DSLs. Furthermore, we capture the availability of attacker models and defense mechanisms.
First, we distinguished between problem and solution oriented DSLs. Most languages (74 %) focus on solutions, e.g., by enhancing defense mechanism implementation and 23 % of the DSLs focus on (identifying) problems. A few DSLs (3 %) are both, problem and solution oriented.
Security objectives
We investigated which security objectives the security DSLs address and were able to map the DSLs to seven fundamental security objectives (cf. Fig. 4(a)).
A DSL can address multiple security objectives, some DSLs are not tied to a specific objective and can be used for any potential security objective. This can be the case, for example, for threat modeling languages. However, the security objectives are explicitly formulated for only 46 % of the security DSLs, sometimes by making them part of the security DSL or the framework around. The other 54 % of the DSLs address security objectives implicitly, e.g., in terms of considered countermeasures.
Most DSLs target one of two core security objectives: confidentiality and integrity. The former is addressed by 55 DSLs (46 %) and the latter by 37 DSLs (31 %). Availability, the third core security objective is only targeted by 13 DSLs (11 %) and derived security objectives like privacy are focused by even fewer languages. However, 48 % of the DSLs do not consider a fixed set of objectives but can potentially target any security objective.
@@
Listing 1: An example in the Paragon language that demonstrates legal and illegal flow of data by direct and indirect flow including the feature of declassification (from Broberg et al. (Broberg et al., 2013a)).
@@
0055551010101015151515202020202525252530303030353535354040404045454545RequirementsAccess ControlThreat ModelingIntrusion DetectionInformation Flow ControlCryptographyNumber of DSLs
@@
Figure 4. Security aspects addressed by the security DSLs
Categories of security concerns
Furthermore, we investigated which security concerns the different DSLs address. We identified six different categories of targeted security concerns (cf. Fig. 4(b)). Thereby, a single security DSL can target multiple of the identified categories.
@@
Of the 120 DSLs we discovered, 23 focus on security requirements. Some of them are tied to business process models, while others standardize the specification of requirements with policies, for example for web services. A typical example for this category is SecureTROPOS (Mouratidis and Giorgini, 2007), a visual, graph-based DSL aimed at the identification and validation of security requirements for multiagent systems.
@@
Access control is the most dominant category we identified (43 DSLs). Multiple DSLs of this category focus on effective specification of access control policies. Furthermore, languages in this category are used to define access control models, such as the UML-based access control modeling language (UACML(Slimani et al., 2011).
@@
We found 17 DSLs that can be placed in the threat modeling category. Some of them facilitate the identification of attack vectors based on design models. Others allow efficient descriptions of attacks for simulations and to detect potential countermeasures. Part of this category is, e.g., MAL (Johnson et al., 2018), a probabilistic meta threat modeling and attack simulation language
@@
DSLs for intrusion detection systems are used to specify known attack patterns which can then be identified during runtime by the intrusion detection system and yield a response. For example, Listing LABEL:lst:statl demonstrates how the fundamental concepts of states and their transitions in STATL (Eckmann et al., 2002) allow expressing potentially vulnerable network protocol states to detect attacks. We mapped 9 DSLs to this category.
@@
Languages for information flow control are typically equipped with a type system that features dedicated types to address confidentiality (secret vs. public) and integrity (trusted vs. untrusted). Via type inference, the flow of sensitive information over variables is automatically tracked, and the framework ensures dynamically or statically that, e.g., secret variables do not reach publicly observable outputs. We mapped 23 DSLs to this category; a well known representative is Jif (Pullicino, 2014).
@@
Cryptographic primitives are usually described with a mixture of continuous text, pseudocode and mathematical expressions, leading to ambiguous specifications that are not suitable to formally verify the equivalence of their implementations. This category includes DSLs tackling this problem by including language constructs that enable an explicit and efficient expression of the underlying primitives (Cryptol (Erkök and Matthews, 2009)). LABEL:lst:cryptol shows how a part of the AES encryption can be specified with Cryptol using its built-in type for polynomials, primitive polynomial modulus, and multiplication operations. Other DSLs in this category enable a simple and correct usage of cryptography, or focus on side-channel secure implementations of cryptography (FaCT (Cauligi et al., 2019)). Some DSLs allow automated generation of MPC or FHE protocols from the function that is computed on the sensitive data (Wysteria (Rastogi et al., 2014)) and thus allows non-cryptographic experts to use this advanced cryptography. In total, 19 DSLs focus on cryptography.
Listing 2: State transition in STATL to identify half-open TCP connections (from Eckmann et al. (Eckmann et al., 2002)).
@@
Listing 3: Specification of multiplication in the Galois Field including the irreducible polynomial for AES written in Cryptol (from Cryptol (Erkök and Matthews, 2009))
@@
Our categorization covers a wide range of subdomains within information security and indicates that DSLs have already been proposed and developed for most relevant security aspects. It is noteworthy that most categories discussed above represent some type of defense mechanism (Access Control, Intrusion Detection, Information Flow Control, and Cryptography), thus potentially indicating that security DSLs may focus on concrete defense mechanisms. In fact, 82 % of the DSLs support a concrete defense mechanism. Many languages that are not associated with a specific defense mechanism aim to specify security requirements, and the choice of defense mechanisms is left to the user. Nevertheless, defenses play an important role in the use of these DSLs as well.
System Scope
Security concerns and defenses are applied with different scopes of the information system. For example, threat modeling can be performed using the same principles at a high level, where the entire enterprise is in scope, and at a lower level, where only the software is considered. Different system scopes present unique challenges, for example, quantifying probabilities for human behavior is likely to be more difficult than quantifying probabilities of success for brute force attacks. We used a coarse-grained categorization of information system scopes that reflects the common structuring of security research domains. We mapped each DSL to a single scope that most closely matches the target scope of its security concept or countermeasure.
@@
We have sorted 73 DSLs (61 %) into the Information Technology (IT) system scope. In this category multiple subsystems combined over, e.g., a network are relevant, but also the interaction of multiple programs, including operating systems.
@@
Twenty DSLs (17 %) focus on the security of individual software, for example, software implementations of cryptographic primitives.
@@
Although this work aims at DSLs for software-intensive systems, we also included 8 (7 %) languages that focus on the security of individual hardware components. These are primarily (programming-language like) hardware description languages.
@@
Many security DSLs (36 %) come with an explicitly specified attacker model. The categories cryptography and information flow control stand out with 87 % and 84 %, respectively.
@@
For a number of DSLs (15 %) it is not possible to specify a concrete attacker model since it depends on the concrete subject system to which this DSL is applied. These DSLs are mostly design-time security DSLs.
@@
There are also some security DSLs (7 % ) that do not come with an attacker model since it is their purpose to allow security experts to specify such an attacker model, e.g., as part of threat modeling.
@@
Unfortunately, with 51 security DSLs (43 %), a significant number does not provide an attacker model. While for some security DSLs it is not feasible to specify an attacker model, e.g., DSLs for reasoning about security requirements, these are only a few exceptions of the DSLs not providing an attacker model.
@@
4.3. RQ3: What phases of the development process are supported by the DSLs?
@@
Relevant phases of the SDLC
The DSLs are used in six different phases of the SDLC, whereby 44 % of the DSLs are relevant in multiple phases. On average, every DSL is used in 1.5 phases. Figure 5 shows how the DSLs distribute across the phases of the SDLC and concrete numbers of how often each phase is relevant for a security DSL, in which phases instances of the DSLs are created, and in which phases these are actually put into use.
@@
Req.DesignCodeTestDepl.Runtime0020202020404040406060606080808080100100100100Number of DSLsinstantiationrelevantuse
@@
Figure 5. Relevant SDLC Phases for the Usage of Security DSLs
@@
After the requirements phase, 49 DSLs (41 %) are instantiated in the design phase. Another six DSLs instantiated in the requirements phase are relevant to the design phase (55 DSLs in total). In the design phase, more-detailed security requirements are specified in relation to the system’s design. Here, a wide range of design models with different levels of abstraction are supported. For example, UMLsec allows to explicitly classify information such as sensitive information in abstract domain models and at the same time allows planning detailed requirements on communication security, e.g., whether a connection needs to be encrypted, on detailed deployment diagrams (see Fig. 8). Design-level security DSLs are mainly used in combination with the requirements phase and concrete code. Consequently, 40 DSLs are put into use in this phase. However, their use is also relevant in the test phase and at runtime.
@@
With 76 DSLs, the majority of all DSLs (63 %) addresses the implementation phase on source code. Of these DSLs, 67 DSLs are also instantiated in this phase. Code-only DSLs usually provide the means to implement secure code, e.g., containing security labels for secure data flow analysis, or for securely implementing cryptographic code. While most DSLs are only relevant to the code phase, there is still a significant number that also addresses design-time or runtime, or all three of these phases. We observed no DSL that is used in combination with any other phase except unless it also covers the design-phase. Mostly, concrete security objectives are planned at design-time, implemented or verified in code, thereby putting 54 of the relevant DSLs into active use in the implementation phase, and probably also enforced at runtime.
@@
The fewest DSLs, only 2 DSLs (2 %), are used in the deployment phase. Since DSLs are used in the test phase, they are also used in other phases, such as planning security goals for virtualized environments and actually deploying them securely.
@@
Finally, 52 DSLs are relevant to the runtime of the system, often however the DSL instance is not actively used but transformed into another format before. Only 26 DSLs (22 %) are immediately used at runtime, meaning that the DSL instance is actively used at runtime. For security DSLs at runtime, we can observe two different kinds of usage. First, security requirements specified at design-time being monitored at runtime, and second, concrete monitoring rules implemented in models or code being executed at runtime.
Figure 6. Partial actor diagram for an electronic health care system in STS.
Artifacts specified using the DSLs
Following the development phases in which the security DSLs are used, we observed a wide range of artifacts that are expressed using the DSLs (cf. Figure 7).
[…]

### v3

@@ 4.1.1. Background of the DSLs
Most papers (49 %) are published in venues targeting the security community, followed by 11 % of venues in software engineering. Other venues belong to the networks (9 %), applied computing (7 %), information systems (7 %), programming languages (7 %), and distributed computing (6 %) communities. Note that a single venue may target multiple communities. We could not assign a significant portion of the venues (3 %) to specific communities, because they target computer science in general. In Fig. 2(b), we list all conferences and journals where at least four of the papers we examined were published, as well as the assigned communities.
@@
01010202030304040505060607070SecuritySoftware EngineeringNetworkProgramming LanguagesInformation SystemsApplied ComputingDistributed ComputingTheoretical Computer ScienceOperating SystemsHardwareRequirementsArtificial IntelligenceOtherNumber of Papers
@@
Only 40 of the 120 DSLs (33 %) have a publicly accessible repository or website. Note that our initial search for publications covers the time span between 2010 and 2022, and with the backwards snowballing we further identified relevant papers dating back until 1997. Many DSLs have been updated even after their release, as shown in Fig. 3. In particular, at the beginning of 2025, 25 % of the updates happened recently in 2023 (2 DSLs) or 2024 (8 DSLs), indicating that quite a few of them are actively maintained and possibly also used in practice or academia.
1997199719981998199919992000200020012001200220022003200320042004200520052006200620072007200820082009200920102010201120112012201220132013201420142015201520162016201720172018201820192019202020202021202120222022202320232024202402244668810101212Number of DSLspublishedlast update
@@
Observation 1. Security DSLs were primarily published in the mid-2010s, mainly from the security community, but also from other communities, such as software engineering. While few DSLs are publicly available, many of them are still actively maintained.
@@ 4.2. RQ2: What Security Aspects are Targeted by the Security DSLs?
To obtain an over. over the security aspects covered by security DSLs, we analyze their fundamental security objectives and provide a coarse-grained categorization according to the security concept and the system scope targeted by each DSLs. Furthermore, we capture the availability of attacker models and defense mechanisms.
First, we distinguished between problem- and solution-oriented DSLs. Most languages (74 %) focus on solutions, e.g., by enhancing defense mechanism implementation and 23 % of the DSLs focus on (identifying) problems. A few DSLs (3 %) are both, problem and solution oriented.
4.2.1. Security Objectives
We investigated which security objectives the DSLs address and were able to map the DSLs to seven fundamental security objectives (cf. Fig. 4(a)).
A DSL can address multiple security objectives, while some DSLs are not tied to a specific objective and can be used for any potential security objective. This can be the case, for example, for threat modeling languages. However, the security objectives are explicitly formulated for only 46 % of the DSLs, sometimes by making them part of the DSL or the framework around. The other 54 % of the DSLs address security objectives implicitly, e.g., in terms of considered countermeasures.
Most DSLs target one of two core security objectives: confidentiality and integrity. The former is addressed by 55 DSLs (46 %) and the latter by 37 DSLs (31 %). Availability is only targeted by 13 DSLs (11 %) and objectives like privacy are focused by even fewer languages. However, 48 % of the DSLs do not consider a fixed set of objectives, but can potentially target any security objective.
@@ 4.2.1. Security Objectives
Listing 1: An example in the Paragon language that demonstrates legal and illegal flow of data by direct and indirect flow including the feature of declassification (from Broberg et al. (242)).
@@
05510101515202025253030353540404545RequirementsAccess ControlThreat ModelingIntrusion DetectionInformation Flow ControlCryptographyNumber of DSLs
@@
Figure 4. Security aspects addressed by the DSLs
4.2.2. Categories of Security Concerns
Furthermore, we investigated which security concerns the different DSLs address. We identified six categories of targeted security concerns (see Fig. 4(b)). Thereby, a single DSL can target multiple of the identified categories.
@@ 4.2.2. Categories of Security Concerns
23 DSLs focus on security requirements. Some of them are tied to business process models, while others standardize the specification of requirements with policies, e.g., for web services. A typical example for this category is SecureTROPOS(211), a visual, graph-based DSL aimed at the identification and validation of security requirements for multiagent systems.
@@
Access control is the most dominant category we identified (43 DSLs). Multiple DSLs of this category focus on effective specification of access control policies. Furthermore, languages in this category are used to define access control models, such as the UML-based access control modeling language (UACML) (71).
@@
We found 17 DSLs that can be placed in the threat modeling category. Some of them facilitate the identification of attack vectors based on design models. Others allow efficient descriptions of attacks for simulations and to detect potential countermeasures. Part of this category is, e.g., MAL(22), a probabilistic meta threat modeling and attack simulation language.
@@
DSLs for intrusion detection systems (9 DSLs) are used to specify known attack patterns which can then be identified at runtime by the intrusion detection system and yield a response. Listing LABEL:lst:statl demonstrates how STATL (209) uses the fundamental concepts of states and transitions to express potentially vulnerable network protocol states for attack detection.
@@
Languages for information flow control (23 DSLs) are typically equipped with a type system that features dedicated types to address confidentiality (secret vs. public) and integrity (trusted vs. untrusted). Via type inference, the flow of sensitive information over variables is automatically tracked, and the framework ensures dynamically or statically that, e.g., secret variables do not reach public outputs. A well known representative is Jif(188).
@@
Cryptographic primitives are usually described with a mixture of plain text, pseudocode, and mathematical expressions, leading to ambiguous specifications. This category includes DSLs tackling this problem by including language constructs that enable an explicit and efficient expression of the underlying primitives (Cryptol(309)). LABEL:lst:cryptol shows how a part of the AES encryption can be specified with Cryptol using its built-in type for polynomials, primitive polynomial modulus, and multiplication operations. Other DSLs in this category enable a simple and correct usage of cryptography, or focus on side-channel secure implementations (FaCT(197)). Some DSLs allow automated generation of MPC or FHE protocols from the specification of a function to be computed on the sensitive data (Wysteria(37)), thus allowing non-cryptography experts to use this advanced cryptography. In total, 19 DSLs focus on cryptography.
Listing 2: State transition in STATL to identify half-open TCP connections (from Eckmann et al. (209)).
@@
Listing 3: Specification of multiplication in the Galois Field including the irreducible polynomial for AES written in Cryptol (from Cryptol (309))
@@
Our categorization covers a wide range of subdomains within information security and indicates that DSLs have already been proposed and developed for most relevant security aspects. It is noteworthy that most categories discussed above represent some type of defense mechanism, indicating that security DSLs focus on defense mechanisms. In fact, 82 % of the DSLs support a concrete defense mechanism. Many languages that are not associated with a specific defense mechanism aim to specify security requirements, and the choice of defense mechanisms is left to the user. Nevertheless, defenses play an important role in the use of these DSLs as well.
4.2.3. System Scope
Security concerns and defenses are applied considering different scopes of the information system. For example, threat modeling can be performed using the same principles at a high level, where the entire enterprise is in scope, and at a lower level, where only the software is considered. We used a coarse-grained categorization of information system scopes that reflects the common structuring of security research domains. We mapped each DSL to a single scope that most closely matches the target scope of its security concept or countermeasure.
@@ 4.2.3. System Scope
We have assigned 73 DSLs (61 %) to the Information Technology (IT) system scope. This category contains combinations of multiple subsystems, e.g., via a network, but also the interaction of multiple programs, including operating systems.
@@
20 DSLs (17 %) focus on the security of individual software, for example, software implementations of cryptographic primitives.
@@
8 DSLs (7 %) focus on the security of individual hardware components. These are primarily (programming-language like) hardware description languages.
@@ 4.2.4. Attacker Model
Many DSLs (36 %) come with an explicitly specified attacker model. The categories cryptography and information flow control stand out with 87 % and 84 %, respectively.
@@
For a number of DSLs (15 %) it is not possible to specify a general attacker model since it depends on the concrete subject system to which this DSL is applied. These DSLs are mostly design-time DSLs.
@@
There are also some DSLs (7 % ) that do not come with an attacker model since it is their purpose to specify such an attacker model, e.g., as part of threat modeling.
@@
Unfortunately, with 51 DSLs (43 %), a significant number does not provide an attacker model. While for some DSLs it is not feasible to specify an attacker model, e.g., DSLs for reasoning about security requirements, these are only a few among the DSLs not providing an attacker model.
@@
4.3. RQ3: What Phases of the Development Process are Supported by the Security DSLs?
@@
4.3.1. Relevant Phases of the SDLC
The DSLs are used in six different phases of the SDLC, whereby 44 % of the DSLs are relevant in multiple phases. On average, every DSL is used in 1.5 phases. Figure 5 shows how the DSLs distribute across the phases of the SDLC and concrete numbers of how often each phase is relevant for a DSL, in which phases instances of the DSLs are created, and in which phases these are actually put into use.
@@ 4.3.1. Relevant Phases of the SDLC
Req.DesignCodeTestDepl.Runtime02020404060608080100100Number of DSLsinstantiationrelevantuse
@@
Figure 5. SDLC phases relevant for using the security DSLs
@@
49 DSLs (41 %) are instantiated in the design phase. Another six DSLs instantiated in the requirements phase are relevant to the design phase. In the design phase, more-detailed security requirements are specified. Here, a wide range of design models with different levels of abstraction are supported. For example, UMLsec allows to explicitly classify information in abstract domain models and allows planning detailed requirements on communication security, e.g., whether a connection needs to be encrypted(see Fig. 8). Design-level DSLs are mainly used in combination with the requirements phase and concrete code. Consequently, 40 DSLs are put into use in this phase. However, their use is also relevant in the test phase and at runtime.
@@
With 76 DSLs, the majority of all DSLs (63 %) addresses the code phase on source code. Of these DSLs, 67 DSLs are also instantiated in this phase. Code-only DSLs usually provide the means to implement secure code, e.g., containing security labels for secure data flow analysis, or for securely implementing cryptographic code. While most DSLs are only relevant to the code phase, there is still a significant number that also addresses design-time or runtime. Except for DSLs used at runtime, we observed no code-level DSL that is used in combination with any other phase unless it also covers the design phase. Mostly, concrete security objectives are planned at design-time, implemented or verified in code, thereby putting 54 of the relevant DSLs into active use in the implementation phase, and probably also enforced at runtime.
@@
The fewest DSLs, only 2 DSLs (2 %), are used in the deployment phase. These DSLs are used in other phases, such as planning security goals for virtualized environments, and also support the actual secure deployment of the system.
@@
Finally, 52 DSLs are relevant to the runtime of the system, but often the DSL instance is not actively used, but is transformed into another format beforehand. For only 26 DSLs (22 %) the DSL instance is actively used at runtime. For DSLs at runtime, we can observe two different kinds of usage. First, security requirements specified at design-time being monitored at runtime, and second, concrete monitoring rules implemented in models or code being executed at runtime.
Security DSLs primarily focus on the code and design phases. During these phases, the majority of DSLs are intended to be instantiated in order to provide a concrete security benefit. Besides these two phases, security DSL instances are often used at runtime. Security DSLs seem to barely address the test and deployment phases.
Figure 6. Partial actor diagram for an electronic health care system in STS
4.3.2. Artifacts Specified Using the DSLs
Following the development phases in which the DSLs are used, we observed a wide range of artifacts that are expressed using security DSLs (cf. Figure 7).
Policy36 % Code27 % Design Model13 % Attack Description9 % Requirements7 % HW Description4 % Intrusion Detection Rule3 % Process Model2 % Security Configuration